# Complementary Probes of Warped Extra Dimension: Colliders, Gravitational Waves and Primordial Black Holes from Phase Transitions

### v1

@@ IV Primordial black hole formation
In the right panel of Fig. 3 we exhibit contour lines of the PBH initial spin. Its value is mild, typically $\sim 10^{-3}$. However, there exist various mechanisms that can enhance the initial spin of PBHs. For example, PBHs can undergo accretion and gain spin De Luca et al. (2020), whereas PBH mergers Hofmann et al. (2016), or close hyperbolic encounters Jaraba and Garcia-Bellido (2021), can do the job in convenient astrophysical environments. Moreover, in the presence of beyond the Standard Model localized at the TeV brane, PBHs can spin up e.g. through the emission of light axion or axion-like particles (which could be measured within 10% by future gamma-ray observatories, provided that $S_{\rm PBH}>0.1$Calzà et al. (2024a, b). PBHs with spin also exhibit what is known as superradiance Yang et al. (2024); Tsukada et al. (2019); Berti et al. (2019); Arvanitaki and Dubovsky (2011), which could also lead to observable signatures, and modify the initial spin of PBH. Spinning PBHs also may show as primordial features in the SGWB spectrum as investigated, originating from second-order tensors Bhaumik et al. (2023) and cosmic strings in the early universe Ghoshal et al. (2023). Depending on the specific mechanism, the final spin can be enhanced up to two orders of magnitude, from the initial spin of PBHs created. Moreover, a method for measuring the PBH spin using the peak amplitude of the GW strain has been described in Ferguson et al. (2019). For a recent study of PBH spin formation in arbitrary FOPTs, see Banerjee and Dey (2024); Banerjee and Harada (2024).
V Gravitational waves
@@
References Megías et al. (2020, 2023) detail the FOPT SGWB predictions for the specific high- and low-energy setups of Sec. II. Here we briefly revisit those computations in view of the new semi-analytic approximations provided in Sec. III and the updates on the SGWB frequency shape recently summarized in Ref. Caprini et al. (2024).
@@
As motivated in Ref. Megías et al. (2018), the FOPTs in our setups are so supercooled ($\alpha\gtrsim 10^{7}$) that one expects the bubble to expand at the speed of light and the plasma to play a negligible role in the GW production. In this regime where the GW signal is mainly sourced by the bubble collision mechanism, the SGWB critical energy density per logarithmic frequency interval is given by Caprini et al. (2024)
@@
The SGWB spectrum then exhibits a broken-power-law frequency shape. In particular, at $f\gg f_{p}$, it behaves as $\Omega_{\textrm{GW}}(f)\propto f^{-2.4}$ and is much steeper than what Refs. Megías et al. (2020, 2023) considered (based on the now outdated Ref. Caprini et al. (2020)), i.e. $\Omega_{\textrm{GW}}(f)\propto f^{-1}$.
We compute $h^{2}\bar{\Omega}_{\textrm{GW}}$ and $f_{p}$ arising in our benchmark points P1-P8. Table 2 quotes the results while Figure 4 displays the corresponding SGWB spectra. Note that, in every pair of benchmark points, P1-P2, P3-P4, $\dots$, the only different input is $N$, being either 10 or 50 (and a small adjustment in $\beta/H$, c.f. Table 2); keeping the same inputs and varying $N$ in the range $5010<N<$ 50 would then move the SGWB signal inside the colored strips in the figure. As expected from Eq. (29), $h^{2}\bar{\Omega}_{\textrm{GW}}$ is essentially constant throughout our benchmark points where $\beta/H$ little varies (and the dependence on $\alpha$ disappears as $\alpha\gg 10$). For the same reason, $f_{p}$ is mainly controlled by $T_{R}$ whose dependence on $N$, $\rho$ and $\beta/H$ is discussed in Sec. II. We see in Fig. 5 (left panel) how such a dependence impacts $f_{p}$ when $\beta/H$ is varied to maximize the PBH abundance bound that is experimentally allowed (c.f. gray border in the right panel of Fig. 2).
Experiments can probe a large portion of the parameter space $\{h^{2}\bar{\Omega}_{\textrm{GW}},f_{p}\}$ between now and the far future. The first class of probes leverages the CMB and BBN observations constraining the radiation energy density in the late universe. Their constraint can be cast in terms of $\Delta N_{\rm eff}=N_{\rm eff}-N_{\rm eff}^{\rm SM}$, the number of additional degrees of freedom beyond the SM ones that are relativistic at the time of recombination: $\Omega_{N_{\rm eff}}\lesssim 5.6\times 10^{-6}\;\Delta N_{\text{eff}}$, with $N_{\rm eff}^{\rm SM}=3.0440(2)$ Akita and Yamaguchi (2020); Froustey et al. (2020); Bennett et al. (2021). Crucially, as for the energy budget of the universe, the SGWB acts as an extra contribution to the radiation energy density, and the CMB and BBN bound holds. For the SGWB spectrum in Eq. (28), this bound implies an upper limit on $h^{2}\bar{\Omega}_{\textrm{GW}}$:
@@
where $f_{\text{min}}\sim 10^{-10}\,\text{Hz}$ for the BBN bound and $f_{\text{min}}\sim 10^{-18}\,\text{Hz}$ for the CMB bound, but the integral is numerically insensitive to the specific value of $f_{\text{min}}$ for signals peaking at much higher frequencies. Using the PLANCK 2018 and BBN bound, $N_{\rm eff}=2.99\pm 0.17$ Workman et al. (2022), we obtain $\Delta N_{\rm eff}\equiv N_{\rm eff}-N_{\rm eff}^{\rm SM}<0.279$ at 95% C.L. and, in turn, $h^{2}\bar{\Omega}_{\text{GW}}\leq 7.2\times 10^{-7}$ at 95% C.L. Moreover, future CMB observations at CMB-S4 Abazajian et al. (2022a, b) and CMB-HD Sehgal et al. (2019); Aiola et al. (2022) are estimated to probe deviations equivalent to $\Delta N_{\rm eff}\sim 0.06$ and $\Delta N_{\rm eff}\sim 0.027$ at 95% C.L., i.e. $h^{2}\bar{\Omega}_{\rm GW}<1.5\times 10^{-7}$ and $h^{2}\bar{\Omega}_{\rm GW}<6.8\times 10^{-8}$, respectively. We display these future reaches and bound in Fig. 4 as a horizontal gray band and horizontal solid lines.
@@
The second class of probes employs GW direct detections. Current PTA and LVK measurements fall within this class. The former provides an amplitude range for a SGWB at nanohertz frequencies Afzal et al. (2023); Antoniadis et al. (2023); Reardon et al. (2023); Xu et al. (2023). For illustrative purposes, in Fig. 4 we display the NANOGrav’s violin plot as a PTA representative constraint Afzal et al. (2023) (the violin plots of the other PTA collaboration are similar Agazie et al. (2024)). The interpretation of this signal is still under debate and there exist dedicated analyses aiming at scrutinizing the FOPT origin. Here we limit ourselves to qualitative arguments. First, we see that the radion FOPTs predicted at the benchmark points P1 and P2 would substantially contribute to the observed PTA data. Nevertheless, to fully fit the violin plot, more power is required at $f\simeq 10^{-7}$ Hz. This may (likely) arise from a population of individually-unresolved supermassive BH binaries. Instead, within the radion FOPT picture only, this would require to move away from P1 and P2 by increasing both $h^{2}\bar{\Omega}_{\textrm{GW}}$, and $f_{p}$, i.e. increasing $\rho$ and decreasing $\beta/H_{\ast}$ (or, equivalently, decreasing $|\lambda_{1}|$). This however violates the PBH bound, unless the production mechanism of Sec. IV can be circumvented or made less efficient. On the other hand, radion FOPT with $T_{R}$ slightly below $10^{-4}\,\textrm{TeV}$ are in some tension ($>95\%$ C.L.) with the PTA data for producing an excess of signal in the lowest-frequency violin posteriors; FOPTs with $T_{R}\ll 10^{-4}$ TeV would be as good as those with $T_{R}\gg 10^{-3}$ TeV as their SGWB would be negligible in the PTA frequency band.
@@
The LVK collaboration provides a 95% C.L. upper bound on the SGWB at decihertz frequencies based on the O3 Run Abbott et al. (2021). Their 95% C.L. bound constrains power-law SGWB signals reaching signal-to-noise ratio $\text{SNR}\geq 2$ over one year of data taking, approximately. We recast this bound to a SGWB with the broken-power-law structure in eq. (28) by constructing the SNR $=2$ “peak integrated sensitivity” Schmitz (2021). We thus determine the parameter region in the plane $\{f_{p},h^{2}\bar{\Omega}_{\rm GW}\}$ yielding
@@
where $\tau_{i}$ is the length of the data in seconds, and $h^{2}\Omega_{\text{exp},i}(f)$ is the experiment’s noise curve parametrized as an energy density. [^7] We remind that $\Omega_{\text{exp}}(f)$ is related to the noise power spectral density $S_{h}$ by the equality $\Omega_{\text{exp}}(f)=\left[(2\pi^{2}f^{3})/(3H_{0}^{2})\right]\,S_{h}(f)$ , and to the (dimensionless) noise characteristic strain $h_{c}(f)=\sqrt{fS_{h}}$ by the equality $\Omega_{\text{exp}}(f)=\left[(2\pi^{2}f^{2})/(3H_{0}^{2})\right]\,h_{c}^{2}(f)$. Using $\tau_{\text{LVK O3}}=11\,$months and $h^{2}\Omega_{\text{exp,LVK O3}}$ provided in Ref. Abbott et al. (2021), we obtain the “LVK O3” hatched area in Fig. 4. By construction, any SGWB with the frequency shape in Eq. (28) is ruled out at 95% C.L. if its peak is above the lower border of this area.
@@
With the same procedure, we forecast the FOPT SGWB detection reach of future GW experiments. Specifically, we take the expected noise curves for aLIGO design (from Fig. 1 of Ref. B.P.Abbott et al. (KAGRA(2019)), ET (from Fig. 6 of Ref. Hild et al. (2011), ET-D curve), AION-km (from Fig. 1 in Ref. Badurina et al. (2021)), AEDGE (from Fig. 2 in Ref. Badurina et al. (2021)), BBO (from Fig. 3 of Ref. Yagi and Seto (2011)), DECIGO (from Fig. 3 of Ref. Yagi and Seto (2011)), LISA (from Fig. 1 in Ref. Amaro-Seoane et al. (2017)), $\mu$-ARES (from Fig. 1 in Ref. Sesana et al. (2021)), THEIA (from Eq. (2.3) in Ref. Garcia-Bellido et al. (2021)) and SKA (from Ref. Moore et al. (2015, ) with 100 and 200 millisecond pulsars as input), assume for concreteness $\tau_{\rm aLIGOdes.}=\tau_{\rm ET}=\tau_{\rm AION\,km}=\tau_{\rm AEDGE}=\tau_{\rm LISA}=\tau_{\rm\mu-ARES}=4\,$yr and $\tau_{\rm THEIA}=\tau_{\rm SKA}=20\,$yr, and finally compute the regions of the plane $\{f_{p},h^{2}\bar{\Omega}_{\rm GW}\}$ satisfying the condition in Eq. (32). The lower borders of these regions are presented in Fig. 4, and a FOPT with frequency shape as in Eq. (28) with the peak above a given border would have SNR > $>\,2$ in the corresponding detector. The synergy and complementarity of the considered GW experiments is manifest in Fig. 4.
@@
It is worth stressing that a SGWB source with a given SNR in a detector implies that its signal is present in the data with some statistical significance. Nevertheless, the SNR criterion does not guarantee detection. Foregrounds due to unresolved individual sources may hinder the detection of even SGWBs with high SNR. Moreover, for signal-dominated experiments where the noise curve must be determined together with the signals, the SNR evaluation does not include the uncertainties on the noise reconstruction. To be precise on the detection capabilities, one should run the data analysis pipeline for the SGWB reconstruction data analysis pipeline dedicated to every experiment (see e.g. Refs. Badger et al. (2023); Gowling et al. (2023); Hindmarsh et al. (2024); Caprini et al. (2024) for proof-of-principle examples). Given the complexity of such a task, we limit ourselves to the SNR estimate and let the scientific collaborations reach more quantitative conclusions.
@@
Finally, in the right panel of Fig. 5 we remap the current bounds and forthcoming probes as functions of $T_{R}$ and $\beta/H$. To show the GW experiment synergy expected in the late 40s, we focus on the detection capabilities (based on the simplistic proxy SNR > $>\,2$) of ET, aLIGO design, LISA and SKA with $\tau_{\textrm{aLIGO des.}}=\tau_{\textrm{ET}}=7\,$yr, $\tau_{\textrm{LISA}}=4\,$yr and $\tau_{\textrm{SKA}}=20\,$yr. It results that the future network of GW experiments has a huge indirect discovery power on PBHs, as their sensitivities reach radion FOPTs that lead to PBH abundances of only $f_{\rm PBH}\gtrsim 5\times 10^{-6}$. In particular, the region where PTA data likely require a SGWB is displayed in the light gray and dark gray areas corresponding respectively to Ref. Afzal et al. (2023) and erratum. [^8] We display the NANOGrav result for concreteness. Other PTA experiments obtain essentially the same result Agazie et al. (2024). We see that the benchmark point P1 is inside this region, so the values of the (low-energy) 5D setup parameters favored by the PTA measurement should be around those of P1, in agreement with Ref. Megías et al. (2023). In this remapping, the current BBN and Planck constraint displayed in Fig. 4 rules out the region $\beta/H_{\ast}>1.1$ at 95% C.L., while the future CMB-S4 and CMB-HD observations will further test the region $2.31.1<\beta/H_{\ast}<$ 2.3 and $3.41.1<\beta/H_{\ast}<$ 3.4 at 95% C.L..
Figure 5: Left panel: Contour plot of the peak frequency $f_{p}$ in the plane $\{\rho,N\}$ for $\beta/H_{\ast}=(\beta/H_{\ast})_{\textrm{min}}$, i.e. the minimal value allowed by the PBH constraints in Fig. 3. Right panel: The parameter reach of the present and future GW experiment network. In each area the corresponding experiment (see labels) detect the FOPT SGWB with SNR $>2$. The area of NANOGrav 15yr shows the 95% C.L. favored region of Ref. Afzal et al. (2023), being the lighter area the region shown in the original reference, while the darker area corresponds to the one shown in the erratum of that reference. The BBN bound (horizontal gray band) rules out the region $\beta/H_{\ast}<1.1$. The position of the benchmark points of Table 2 are shown as black bullet points. In the regions labeled as LISA, ET, LVK O3 and LVK O5, the corresponding experiments detect the FOPT SGWB with SNR $>2$.
VI Collider and astrophysical data

### v2

@@ 4 Primordial black hole formation
In the upper right panel of Fig. 3 we exhibit contour lines of the PBH initial spin. Its value is mild, typically $\sim 10^{-3}$. However, there exist various mechanisms that can enhance the initial spin of PBHs. For example, PBHs can undergo accretion and gain spin DeLuca:2020bjf , whereas PBH mergers Hofmann:2016yih , or close hyperbolic encounters Jaraba:2021ces , can do the job in convenient astrophysical environments. Moreover, in the presence of beyond the Standard Model localized at the TeV brane, PBHs can spin up e.g. through the emission of light axion or axion-like particles (which could be measured within 10% by future gamma-ray observatories, provided that $S_{\rm PBH}>0.1$Calza:2021czr ; Calza:2023rjt . PBHs with spin also exhibit what is known as superradiance Yang:2023aak ; Tsukada:2018mbp ; Berti:2019wnn ; Arvanitaki:2010sy , which could also lead to observable signatures, and modify the initial spin of PBH. Spinning PBHs also may show as primordial features in the SGWB spectrum as investigated, originating from second-order tensors Bhaumik:2022zdd and cosmic strings in the early universe Ghoshal:2023sfa . Depending on the specific mechanism, the final spin can be enhanced up to two orders of magnitude, from the initial spin of PBHs created. Moreover, a method for measuring the PBH spin using the peak amplitude of the GW strain has been described in Ferguson:2019slp . For a recent study of PBH spin formation in arbitrary FOPTs, see Banerjee:2023qya ; Banerjee:2024nkv . [^8] In this work, we restrict ourselves to considering only a monochromatic PBH mass distribution defined by the mass inside the sound horizon, unlike Refs. Lewicki:2024ghw ; Baldes:2023rqv which also investigated extended PBH mass distributions. Moreover, for simplicity we do not consider the fluctuation of the nucleation time of the first bubble, unlike Ref. Lewicki:2024ghw which also takes into account the impact of fluctuation of the nucleation time of the first $j_{c}$ bubbles. Finally we remark that Refs. Flores:2024lng ; Lewicki:2023ioy study PBH formation in purely vacuum-energy dominated patches only, before the beginning of any bubble nucleation. Particularly, Ref. Flores:2024lng singles out this scenario and raises concern that curvature perturbation may get efficiently generated after nucleation of bubbles has started. Presence of such curvature may play a role in modifying the dynamics. In the present work however, we treat late-blooming patches as evolving independently from the background and do not consider the effects of the curvature $K$ that would result from their interaction. Going beyond these assumptions would be interesting but is beyond the scope of the current analysis.
5 Gravitational waves
@@
References Megias:2020vek ; Megias:2023kiy detail the FOPT SGWB predictions for the specific high- and low-energy setups of Sec. 2. Here we briefly revisit those computations in view of the new semi-analytic approximations provided in Sec. 3 and the updates on the SGWB frequency shape recently summarized in Ref. Caprini:2024hue .
@@
As motivated in Ref. Megias:2018sxv , the FOPTs in our setups are so supercooled ($\alpha\gtrsim 10^{7}$) that one expects the bubble to expand at the speed of light and the plasma to play a negligible role in the GW production. In this regime where the GW signal is mainly sourced by the bubble collision mechanism, the SGWB critical energy density per logarithmic frequency interval is given by Caprini:2024hue
@@
The SGWB spectrum then exhibits a broken-power-law frequency shape. In particular, at $f\gg f_{p}$, it behaves as $\Omega_{\textrm{GW}}(f)\propto f^{-2.4}$ and is much steeper than what Refs. Megias:2020vek ; Megias:2023kiy considered (based on the now outdated Ref. Caprini:2019egz ), i.e. $\Omega_{\textrm{GW}}(f)\propto f^{-1}$.
We compute $h^{2}\bar{\Omega}_{\textrm{GW}}$ and $f_{p}$ arising in our benchmark points P1-P8. Table 2 quotes the results while Figure 4 displays the corresponding SGWB spectra. Note that, in every pair of benchmark points, P1-P2, P3-P4, $\dots$, the only different input is $N$, being either 10 or 50 (and a small adjustment in $\beta/H$, c.f. Table 2); keeping the same inputs and varying $N$ in the range $5010<N<$ 50 would then move the SGWB signal inside the colored strips in the figure. As expected from Eq. (29), $h^{2}\bar{\Omega}_{\textrm{GW}}$ is essentially constant throughout our benchmark points where $\beta/H$ little varies (and the dependence on $\alpha$ disappears as $\alpha\gg 10$). For the same reason, $f_{p}$ is mainly controlled by $T_{R}$ whose dependence on $N$, $\rho$ and $\beta/H$ is discussed in Sec. 2. We see in the lower panel of Fig. 5 how such a dependence impacts $f_{p}$ when $\beta/H$ is varied to maximize the PBH abundance that is experimentally allowed at any given PBH mass (c.f. gray border in the right panel of Fig. 2). In the same panel, we display the parameter regions that current and forthcoming PTA experiments and interferometers can probe via FOPT SGWB measurements. Such regions are computed as described hereafter.
Experiments can probe a large portion of the parameter space $\{h^{2}\bar{\Omega}_{\textrm{GW}},f_{p}\}$ between now and the far future. The first class of probes leverages the CMB and BBN observations constraining the radiation energy density in the late universe. Their constraint can be cast in terms of $\Delta N_{\rm eff}=N_{\rm eff}-N_{\rm eff}^{\rm SM}$, the number of additional degrees of freedom beyond the SM ones that are relativistic at the time of recombination: $\Omega_{N_{\rm eff}}\lesssim 5.6\times 10^{-6}\;\Delta N_{\text{eff}}$, with $N_{\rm eff}^{\rm SM}=3.0440(2)$ Akita:2020szl ; Froustey:2020mcq ; Bennett:2020zkv . Crucially, as for the energy budget of the universe, the SGWB acts as an extra contribution to the radiation energy density, and the CMB and BBN bounds hold. For the SGWB spectrum in Eq. (28), this bound implies an upper limit on $h^{2}\bar{\Omega}_{\textrm{GW}}$:
@@
where $f_{\text{min}}\sim 10^{-10}\,\text{Hz}$ for the BBN bound and $f_{\text{min}}\sim 10^{-18}\,\text{Hz}$ for the CMB bound, but the integral is numerically insensitive to the specific value of $f_{\text{min}}$ for signals peaking at much higher frequencies. Using the PLANCK 2018 and BBN bound, $N_{\rm eff}=2.99\pm 0.17$ ParticleDataGroup:2022pth , we obtain $\Delta N_{\rm eff}\equiv N_{\rm eff}-N_{\rm eff}^{\rm SM}<0.279$ at 95% C.L. and, in turn, $h^{2}\bar{\Omega}_{\text{GW}}\leq 7.2\times 10^{-7}$ at 95% C.L. Moreover, future CMB observations at CMB-S4 CMB-S4:2020lpa ; CMB-S4:2022ght and CMB-HD Sehgal:2019ewc ; CMB-HD:2022bsz are estimated to probe deviations equivalent to $\Delta N_{\rm eff}\sim 0.06$ and $\Delta N_{\rm eff}\sim 0.027$ at 95% C.L., i.e. $h^{2}\bar{\Omega}_{\rm GW}<1.5\times 10^{-7}$ and $h^{2}\bar{\Omega}_{\rm GW}<6.8\times 10^{-8}$, respectively. We display these future reaches and bounds in Fig. 4 as a horizontal gray band and horizontal solid lines.
Figure 5: Upper panel: The parameter reach of the present and future GW experiment network. In each area the corresponding experiment (see labels) detects the FOPT SGWB with SNR $>2$. The area of NANOGrav 15yr shows the 95% C.L. favored region of Ref. NANOGrav:2023hvm , being the lighter area the region shown in the original reference, while the darker area corresponds to the one shown in the erratum of that reference. The BBN bound (horizontal gray band) rules out the region $\beta/H_{\ast}<1.1$. The position of the benchmark points of Table 2 are shown as black bullet points. In the regions labeled as LISA, ET, LVK O3 and LVK O5, the corresponding experiments detect the FOPT SGWB with SNR $>2$. Lower panel: Contour regions of the peak frequency $f_{p}$ in the plane $\{\rho,N\}$ for $\beta/H_{\ast}=(\beta/H_{\ast})_{\textrm{min}}$, i.e. the minimal value allowed by the PBH constraints in Fig. 2. The parameter reach for SGWB detection at LVK O5, LISA, ET and NANOGrav, along with the excluded region by LVK O3, is derived from the upper panel.
@@
The second class of probes employs GW direct detections. Current PTA and LVK measurements fall within this class. The former provides an amplitude range for a SGWB at nanohertz frequencies NANOGrav:2023hvm ; EPTA:2023fyk ; Reardon:2023gzh ; Xu:2023wog . For illustrative purposes, in Fig. 4 we display the NANOGrav’s violin plot as a PTA representative constraint NANOGrav:2023hvm (the violin plots of the other PTA collaborations are similar InternationalPulsarTimingArray:2023mzf ). The interpretation of this signal is still under debate and there exist dedicated analyses aiming at scrutinizing the FOPT origin. Here we limit ourselves to qualitative arguments. First, we see that the radion FOPTs predicted at the benchmark points P1 and P2 would substantially contribute to the observed PTA data. Nevertheless, to fully fit the violin plot, more power is required at $f\simeq 10^{-7}$ Hz. This may (likely) arise from a population of individually-unresolved supermassive BH binaries. Instead, within the radion FOPT picture only, this would require to move away from P1 and P2 by increasing both $h^{2}\bar{\Omega}_{\textrm{GW}}$, and $f_{p}$, i.e. increasing $\rho$ and decreasing $\beta/H_{\ast}$ (or, equivalently, decreasing $|\lambda_{1}|$). This however violates the PBH bound, unless the production mechanism of Sec. 4 can be circumvented or made less efficient (see e.g. Ref. Hashino:2025fse ). On the other hand, radion FOPT with $T_{R}$ slightly below $10^{-4}\,\textrm{TeV}$ are in some tension ($>95\%$ C.L.) with the PTA data for producing an excess of signal in the lowest-frequency violin posteriors; FOPTs with $T_{R}\ll 10^{-4}$ TeV would be as good as those with $T_{R}\gg 10^{-3}$ TeV as their SGWB would be negligible in the PTA frequency band.
@@
The LVK collaboration provides a 95% C.L. upper bound on the SGWB at decihertz frequencies based on the O3 Run KAGRA:2021kbb . Their 95% C.L. bound constrains power-law SGWB signals reaching signal-to-noise ratio $\text{SNR}\geq 2$ over one year of data taking, approximately. We recast this bound to a SGWB with the broken-power-law structure in eq. (28) by constructing the SNR $=2$ “peak integrated sensitivity” Schmitz:2020syl . We thus determine the parameter region in the plane $\{f_{p},h^{2}\bar{\Omega}_{\rm GW}\}$ yielding
@@
where $\tau_{i}$ is the length of the data in seconds, and $h^{2}\Omega_{\text{exp},i}(f)$ is the experiment’s noise curve parametrized as an energy density. [^9] We remind that $\Omega_{\text{exp}}(f)$ is related to the noise power spectral density $S_{h}$ by the equality $\Omega_{\text{exp}}(f)=\left[(2\pi^{2}f^{3})/(3H_{0}^{2})\right]\,S_{h}(f)$ , and to the (dimensionless) noise characteristic strain $h_{c}(f)=\sqrt{fS_{h}}$ by the equality $\Omega_{\text{exp}}(f)=\left[(2\pi^{2}f^{2})/(3H_{0}^{2})\right]\,h_{c}^{2}(f)$. Using $\tau_{\text{LVK O3}}=11\,$months and $h^{2}\Omega_{\text{exp,LVK O3}}$ provided in Ref. KAGRA:2021kbb , we obtain the “LVK O3” hatched area in Fig. 4. By construction, any SGWB with the frequency shape in Eq. (28) is ruled out at 95% C.L. if its peak is above the lower border of this area.
@@
With the same procedure, we forecast the FOPT SGWB detection reach of future GW experiments. Specifically, we take the expected noise curves for aLIGO design (from Fig. 1 of Ref. LIGOO5 ), ET (from Fig. 6 of Ref. Hild:2010id , ET-D curve), AION-km (from Fig. 1 in Ref. Badurina:2021rgt ), AEDGE (from Fig. 2 in Ref. Badurina:2021rgt ), BBO (from Fig. 3 of Ref. Yagi:2011wg ), DECIGO (from Fig. 3 of Ref. Yagi:2011wg ), LISA (from Fig. 1 in Ref. LISA:2017pwj ), $\mu$-ARES (from Fig. 1 in Ref. Sesana:2019vho ), THEIA (from Eq. (2.3) in Ref. Garcia-Bellido:2021zgu ) and SKA (from Ref. Moore:2014lga ; MooreWebpage with 100 and 200 millisecond pulsars as input), assume for concreteness $\tau_{\rm aLIGOdes.}=\tau_{\rm ET}=\tau_{\rm AION\,km}=\tau_{\rm AEDGE}=\tau_{\rm LISA}=\tau_{\rm\mu-ARES}=4\,$yr and $\tau_{\rm THEIA}=\tau_{\rm SKA}=20\,$yr, and finally compute the regions of the plane $\{f_{p},h^{2}\bar{\Omega}_{\rm GW}\}$ satisfying the condition in Eq. (32). The lower borders of these regions are presented in Fig. 4, and a FOPT with frequency shape as in Eq. (28) with the peak above a given border would have SNR > $>\,2$ in the corresponding detector. The synergy and complementarity of the considered GW experiments is manifest in Fig. 4.
@@
It is worth stressing that a SGWB source with a given SNR in a detector implies that its signal is present in the data with some statistical significance. Nevertheless, the SNR criterion does not guarantee detection. Foregrounds due to unresolved individual sources may hinder the detection of even SGWBs with high SNR. Moreover, for signal-dominated experiments where the noise curve must be determined together with the signals, the SNR evaluation does not include the uncertainties on the noise reconstruction. To be precise on the detection capabilities, one should run the data analysis pipeline for the SGWB reconstruction data analysis pipeline dedicated to every experiment (see e.g. Refs. Badger:2022nwo ; Gowling:2022pzb ; Hindmarsh:2024ttn ; Caprini:2024hue for proof-of-principle examples). Given the complexity of such a task, we limit ourselves to the SNR estimate and let the scientific collaborations reach more quantitative conclusions.
@@
Finally, in the upper panel of Fig. 5 we remap the current bounds and forthcoming probes as functions of $T_{R}$ and $\beta/H$. To show the GW experiment synergy expected in the late 40s, we focus on the detection capabilities (based on the simplistic proxy SNR > $>\,2$) of ET, aLIGO design, LISA and SKA with $\tau_{\textrm{aLIGO des.}}=\tau_{\textrm{ET}}=7\,$yr, $\tau_{\textrm{LISA}}=4\,$yr and $\tau_{\textrm{SKA}}=20\,$yr. It results that the future network of GW experiments has a huge indirect discovery power on PBHs, as their sensitivities reach radion FOPTs that lead to PBH abundances of only $f_{\rm PBH}\gtrsim 5\times 10^{-6}$. In particular, the region where PTA data likely require a SGWB is displayed in the light gray and dark gray areas corresponding respectively to Ref. NANOGrav:2023hvm and erratum. [^10] We display the NANOGrav result for concreteness. Other PTA experiments obtain essentially the same result InternationalPulsarTimingArray:2023mzf . We see that the benchmark point P1 is inside this region, so the values of the (low-energy) 5D setup parameters favored by the PTA measurement should be around those of P1, in agreement with Ref. Megias:2023kiy . In this remapping, the current BBN and Planck constraint displayed in Fig. 4 rules out the region $\beta/H_{\ast}<1.1$ at 95% C.L., while the future CMB-S4 and CMB-HD observations will further test the region $2.31.1<\beta/H_{\ast}<$ 2.3 and $3.41.1<\beta/H_{\ast}<$ 3.4 at 95% C.L..
6 Collider and astrophysical data